# Conceptualization, Data Curation, Formal Analysis, Methodology, Software, Writing – original draft

### v1

@@ 4 Results and Discussion
This section begins by outlining the evaluation metrics used to assess the model after updating the parameters of the cross-modality adaptation network $f_{\text{xa}}(\cdot)$ and the regression head $f_{h}(\cdot)$. We then compare model performance across different LLM backbones to examine how linguistic capability affects prediction accuracy. Finally, we conduct a prompt structure analysis to evaluate the importance of each type of aeronautical information and its contribution to delay prediction accuracy.
@@ 4.1 Evaluation Metrics
To quantify the model’s predictive accuracy, we first evaluate its ability to predict the post-terminal duration $y_{i}\equiv\Delta t_{post,i}$ directly from the scenario input $s_{i,t}$, formulated as $\hat{y}_{i}=f_{\theta}(s_{i,t})$. Ultimately, we assess the model’s effectiveness in estimating the total delay $D_{i}$, by incorporating the known actual entry time $T^{act}_{entry,i}$ and scheduled arrival time $T^{sch}_{arr,i}$, expressed as $\hat{D}_{i}=\left(T^{\text{act}}_{\text{entry},i}-T^{\text{sch}}_{\text{arr},i}\right)+\hat{y}_{i}$. The performance of both predictions is evaluated using a set of standard regression metrics. These metrics collectively capture various aspects of model performance, including the average magnitude of prediction errors, their squared severity, relative error scaling, and the proportion of variance explained by the model.
  • Mean Absolute Error (MAE) measures the average magnitude of errors in predictions, without considering their direction. The evaluation on total delay is given by:
@@
    $=\frac{1}{N}\sum_{i=1}^{N}\left|\left(T_{entry}^{act}-T_{arr}^{sch}+\hat{y}_{i}\right)-\left(T_{entry}^{act}-T_{arr}^{sch}+y_{i}\right)\right|$ (26)
@@
    Therefore, the evaluation on the total delay is equivalent to evaluating the prediction of the post-terminal duration. MAE is easy to interpret and robust to outliers, as it represents the average magnitude of errors in absolute terms. However, it may not adequately reflect the impact of larger errors due to its linear penalty formulation.
@@
    $=\frac{1}{N}\sum_{i=1}^{N}\left(\left(T_{entry}^{act}-T_{arr}^{sch}+\hat{y}_{i}\right)-\left(T_{entry}^{act}-T_{arr}^{sch}+y_{i}\right)\right)^{2}$ (29)
@@
    Due to the absolute terms in SMAPE’s denominator, we separately evaluate delay prediction ($\text{SMAPE}_{D}$) and post-terminal duration prediction ($\text{SMAPE}_{\Delta t}$). SMAPE is used instead of MAPE to reduce instability from zero ground truth values, common in on-time flights. While SMAPE can still be unstable when true or predicted values are near zero, it is generally more robust than MAPE.
@@
    Due to the presence of the mean in the denominator, the $R^{2}$ score for total delay ($R^{2}_{D}$) cannot be equivalent to that of the post-terminal duration ($R^{2}_{\Delta t}$); thus, they are evaluated separately. The $R^{2}$ score provides an intuitive measure of goodness-of-fit, indicating the proportion of variance explained by the model. While it can be misleading for small or imbalanced datasets, this is less of a concern given the scale and diversity of our data.
4.2 Model Performance
@@
This section illustrates the performance of our framework when integrated with various LLM backbones, allowing for a direct comparison of how the varying linguistic capabilities of pre-trained LLMs with different architectures and training strategies impact prediction accuracy. Our evaluation protocol trains the overall model while replacing the LLM backbone with various candidates. For each backbone, we conduct experiments on 12 monthly datasets from 2022, training and testing separately on each month’s corresponding train-test split, visualized in Figure 3. Each scenario $s_{i,t}$ contains the full context, including flight plan, weather, NOTAMs, and trajectories representing airspace conditions. This full context is provided to the model to indirectly predict the total delay $D_{i}$ through the estimation of the post-terminal duration $\hat{y}_{i}=f_{\theta}(s_{i,t})$, using hyperparameters strictly following the reproduction details.
@@
Table 3 presents a performance comparison of several lightweight LLMs on the flight delay prediction task. Among the models evaluated, LLaMA-3.2-1B-Instruct consistently outperforms others across all key metrics for many months. We hypothesize that its high-quality pretraining corpus contributes to its ability to generalize to aeronautical information. Its instruction-tuned training enhances the ability to follow structured prompts, and despite having relatively few parameters, it remains capable of capturing long context in post-terminal duration prediction.
@@
Across the evaluated models, those trained primarily as English single-language models, unlike the Qwen-based model, tend to outperform multilingual models, suggesting that linguistic alignment with the structured English prompts in our framework provides an advantage. Interestingly, DeepSeek‐R1‐Distill‐Qwen‐1.5B, although distilled from the math‐specialized Qwen‐2.5‐1.5B‐Math model, performs better than its Qwen counterparts, suggesting that the distillation process may have introduced useful generalization capabilities. The TinyLLaMA variants demonstrate competitive performance, although they generally do not achieve the accuracy of LLaMA-3.2 models. In addition, the Pythia‐1B model achieves strong results, in some cases exceeding LLaMA-3.2-1B-Instruct, suggesting that it can be a suitable candidate under appropriate training conditions.
@@
Although all models are provided with the full context and trained until the trajectory embeddings align with the tokens that enable the LLM to predict delays with minimal error, different LLM backbones still exhibit varying levels of performance. This reflects the importance of the models’ linguistic understanding in achieving accurate post-terminal duration prediction. Thus, it is not only the inclusion of the full context that matters, but also the model’s ability to effectively interpret that context in order to achieve the best prediction results.
In addition to assessing the comparative effectiveness of each model, we demonstrate that our proposed framework is compatible with arbitrary LLM backbones, as it consistently integrates trajectory representations with the language modality for accurate delay prediction across different models. This highlights the potential of the proposed cross-modality adaptation technique for application to larger and more capable models.
@@
Our framework achieves high $R^{2}_{D}$ values due to our unique problem formulation, which leverages information already available to ATCs at the destination aerodrome, namely, the scheduled time of arrival and the actual time when the aircraft enters the TMA, marking the start of monitoring. This enables efficient prediction without the need for historical delay data or airport network information, unlike the sequential and graph-based models compared in Table 1. By narrowing the focus to a single airport, the framework aligns with the operational scope of ATCs and achieves significantly more accurate delay prediction, without requiring the ability to monitor many airports simultaneously.
From a practical perspective, our framework with several backbones maintains prediction error within one minute and can be considered deployable, as delays are typically recorded at the minute level [75]. While second-level accuracy may be desirable for certain operational needs, the framework already meets the current standard of minute-level records, and future access to higher-resolution data could provide opportunities for even greater precision.
@@ 4.3 Effectiveness of Context Inclusion
A key contribution of our proposed framework is the integration of a comprehensive set of information, providing a representation of the operational environment to the model, including general flight information, weather reports and forecasts, NOTAMs, and airspace conditions modeled through three types of trajectories set. These contextual components are of varying importance; therefore, to evaluate their contributions, we adopt an ablation-based approach in which each component is selectively removed from the multimodal scenario representation $s_{i,t}$, followed by retraining and re-evaluation under strictly identical hyperparameter settings. We used Meta’s LLaMA-3.2-1B-Instruct as the LLM backbone, for the results presented in Table 3 show that the model consistently outperforms other LLMs across the majority of datasets and evaluation metrics. This controlled procedure enables us to directly quantify the performance degradation resulting from the absence of each context type, thereby providing empirical evidence for its contribution to the overall predictive capability. Moreover, this analysis was conducted to support our design choice of using full context in the final configuration, demonstrating the value of leveraging multimodal aeronautical data for accurate prediction of $\Delta t_{post}$. The importances of context elements are discussed as follows:
@@
  • General Flight Information: The exclusion of basic flight information leads to a clear drop in performance, underscoring the importance of general operation details of the flight. This finding aligns with prior works, which predominantly rely on general flight information as the primary input for delay prediction, as shown in Table 1.
@@
  • NOTAMs: Excluding NOTAMs shortens the contextual input, which can sometimes reduce the burden of long sequences. However, it also removes special operational constraints or disruptions, such as runway closures or equipment failures, which can introduce sources of delay [3]. As a result, performance is noticeably reduced overall. This supports our design choice of incorporating NOTAMs as input for post-terminal duration prediction, thereby improving delay estimation.
@@
  • Trajectory Data: Trajectory data is shown to be the most influential factor for post-terminal duration prediction, as its exclusion results in a severe reduction in performance. This is expected, since trajectory data inherently reflect real-time airspace awareness that cannot be inferred from static flight or weather information alone. From a technical perspective, removing trajectory inputs also eliminates the learnable portion of the model’s context. This suggests that integrating trajectory data contributes to the predictive accuracy of our framework.
@@
    • Focusing Trajectories We find that removing the focusing trajectory leads to a large performance degradation, often comparable to the removal of all trajectory inputs and consistently more severe than excluding other trajectory types. This finding is expected and aligns with [25], as the focusing trajectory embedding captures the specific maneuvering procedures being executed and the state of the current flight under monitoring, which are most directly related to its post-terminal duration.
@@
    • Active Trajectories Excluding the active trajectories from the combined input removes the model’s awareness of the current airspace conditions, particularly the level of traffic congestion and the overall traffic situation at that time. Therefore, the model cannot account for the congestion factor that often extends the post-terminal duration, making accurate prediction more difficult.
@@
    • Prior Trajectories Since prior trajectories provide patterns that may influence the active and focusing trajectories and indirectly reflect ground traffic, excluding them leaves the model unaware of historical flow patterns in both air and ground operations, thereby reducing predictive accuracy.
@@
The experimental results indicate that our accurate predictions are enabled by offering a different formulation from prior works and by incorporating more contextual information within a single airspace, as compared in Table 1. While long contextual inputs can sometimes cause the model to struggle, removing certain elements, such as TAFs and NOTAMs, from many datasets may even improve performance. However, including the full context generally ensures consistent and reliable results, as the model can rely on that information when it is actually required. Since LLMs function as selective sequence encoders with the ability to disregard irrelevant tokens, it is preferable to provide the full context and allow the model to ignore unneeded tokens rather than risk excluding potentially useful information.
@@
The contribution of textual components follows a similar pattern, though the trajectory components play a more critical role. This is partly because trajectory features are trainable and partly because the post-terminal duration is strongly influenced by prevailing traffic conditions. Moreover, durations in other phases of flight are implicitly accounted for through events that occur during the pre-terminal phases. The framework demonstrates robustness even when certain contextual information is unavailable, underscoring its suitability for operational use. Without such a model, air traffic controllers would need to assume an average additional more than 3 minutes for every flight [85], which could introduce unnecessary errors for non-delayed aircraft. In contrast, our model achieves an average error between 0.9214 and 2.7395 minutes, highlighting its accuracy and operational value.
@@
Table 4: Prompt Component Ablation Results using Meta-LLaMA-3.2-1B-Instruct
@@
   Without Flight Information 0.9738 2.2846 1.4955 11.1673 0.99979 7.8250 0.9142
@@
4.4 Demonstration for Second-by-Second Delay Update
@@
Although the model underestimates the delay in this example, we observe that the absolute error generally remains within one minute, with predictions adapting as the flight progresses. Moreover, the variation in error over time suggests that the model refines its predictions as newly available information updates the prediction context. Importantly, maintaining sub-minute error levels is practically useful for ATM, since delays are typically recorded at the minute level [75]; keeping the error below one minute is therefore considered deployable. Thus, the real-time prediction model from our framework can potentially be integrated into post-terminal ATM systems to assist ATCs in estimating the delay of aircraft within the terminal area.
@@ 5 Conclusion
This paper presents a flight delay prediction framework that indirectly estimates post-terminal delays by predicting post-terminal duration. Since flight schedules and airspace entry times are already known to ATCs at the destination airport, the framework is designed to support their operational needs by providing accurate and continuously updated delay information. It employs a cross-modality adaptation technique that aligns multimodal data with LLMs, adapting trajectory embeddings from a pre-trained ATSCC encoder into the language modality. The training process remains lightweight, leveraging the pretrained knowledge of both the trajectory encoder and the LLM, while only updating a cross-modality adaptation network and an output regression head for prediction. The framework enables accurate post-terminal duration prediction, with performance gains driven by the comprehensive inclusion of aeronautical information, especially trajectory data, which contribute the most after the cross-modality adaptation. These inputs allow the model to reference key factors influencing delays, while the linguistic capabilities of LLMs enable the interpretation of complex contextual information. By incorporating a wide range of context within the narrower airspace setting, our framework addresses several key limitations of prior works, such as the reliance on prior delay information, the requirement for structured inputs, and the restriction to coarse update timescales. In contrast, our approach focuses on the real-time situation in the airspace, incorporates prior delays through information already available to ATC, flexibly integrates all necessary context, and supports second-by-second updates, making it particularly well-suited for real-time post-terminal ATM operations. Looking forward, the framework can be extended to other phases of delay prediction by incorporating contextual factors relevant to those phases. Furthermore, its generality opens opportunities for broader ATM tasks, such as trajectory prediction and conflict detection. Future work may also explore the integration of higher-quality operational data and additional modalities, such as images or audio, to further advance multimodal models in aviation research in a practical and easily deployable manner.

### v2

@@ 4 Results and Discussion
This section begins by outlining the evaluation metrics used to assess the model after updating the parameters of the cross-modality adaptation network $f_{\text{xa}}(\cdot)$ and the regression head $f_{h}(\cdot)$. We then compare model performance across different LLM backbones to examine how linguistic capability affects prediction accuracy. We conduct a prompt structure analysis to evaluate the importance of each type of aeronautical information and its contribution to delay prediction accuracy. Finally, we present a demonstration of real-time delay prediction updates upon receiving new information after the aircraft enters the airspace.
@@ 4.1 Evaluation Metrics
To quantify the model’s predictive accuracy, we first evaluate its ability to predict the post-terminal duration $y_{i}\equiv\Delta t_{post,i}$ directly from the scenario input $s_{i,t}$, formulated as $\hat{y}_{i}=f_{\theta}(s_{i,t})$. Ultimately, we assess the model’s effectiveness in estimating the total delay $D_{i}$, by incorporating the known actual entry time $T^{act}_{entry,i}$ and scheduled arrival time $T^{sch}_{arr,i}$, expressed as $\hat{D}_{i}=\left(T^{\text{act}}_{\text{entry},i}-T^{\text{sch}}_{\text{arr},i}\right)+\hat{y}_{i}$. The performance of both predictions is evaluated using a set of standard regression metrics that capture various aspects of model performance.
  • Mean Absolute Error (MAE) measures the average magnitude of errors in predictions, without considering their direction. The evaluation of the total delay is given by:
@@
    $=\frac{1}{N}\sum_{i=1}^{N}\Big|\left(T_{entry}^{act}-T_{arr}^{sch}+\hat{y}_{i}\right)$
    $\quad\quad\quad\quad\quad-\left(T_{entry}^{act}-T_{arr}^{sch}+y_{i}\right)\Big|$ (28)
@@
    Therefore, evaluating the total delay is equivalent to evaluating the prediction of the post-terminal duration. MAE is easy to interpret and robust to outliers, as it represents the average magnitude of errors in absolute terms. However, it may underrepresent the impact of large errors due to its linear penalty formulation.
@@
    $=\frac{1}{N}\sum_{i=1}^{N}\Big(\left(T_{entry}^{act}-T_{arr}^{sch}+\hat{y}_{i}\right)$
    $\quad\quad\quad\quad\quad-\left(T_{entry}^{act}-T_{arr}^{sch}+y_{i}\right)\Big)^{2}$ (31)
@@
    Due to the absolute terms in SMAPE’s denominator, we separately evaluate delay prediction ($\text{SMAPE}_{D}$) and post-terminal duration prediction ($\text{SMAPE}_{\Delta t}$). SMAPE is used instead of MAPE to reduce instability caused by zero ground-truth values, which are common in on-time flights. While SMAPE can still be unstable when true or predicted values are near zero, it is generally more robust than MAPE.
@@
    Because the mean appears in the denominator, the $R^{2}_{D}$ score for total delay ($R^{2}_{D}$) cannot be equivalent to that of the post-terminal duration ($R^{2}_{\Delta t}$); thus, they are evaluated separately. The $R^{2}$ score provides a measure of goodness-of-fit, indicating the proportion of variance explained by the model. While it can be misleading for small or imbalanced datasets, this issue is minor given our dataset’s scale and diversity.
4.2 LLM Backbone Performance
@@
This section evaluates our framework with different LLM backbones to compare how their varying linguistic capabilities affect prediction accuracy. We trained the overall model while replacing the LLM backbone with various candidates. For each backbone, experiments are conducted across our 12 monthly datasets, with separate training and testing on each month’s split, as illustrated in Figure 3. Each scenario $s_{i,t}$ contains the full context, including flight plan, weather, NOTAMs, and trajectories representing airspace conditions. This full context is provided to the model to indirectly predict the total delay $D_{i}$ through the estimation of the post-terminal duration $\hat{y}_{i}=f_{\theta}(s_{i,t})$, using hyperparameters strictly following the reproduction details.
@@
Table 3 and Figure 6 present an average performance comparison of various LLMs on the flight delay prediction. Among the evaluated models, LLaMA-3.2-1B-Instruct consistently outperforms the others across most months and achieves the best overall average performance. We hypothesize that its high-quality pretraining corpus contributes to its ability to generalize to aeronautical information. Its instruction-tuned training enhances the ability to follow structured prompts, and despite having relatively few parameters, it is capable of capturing full context in post-terminal duration prediction.
@@
Across the evaluated models, those trained primarily as English single-language models tend to outperform multilingual models, suggesting that linguistic alignment with the structured English prompts in our framework provides an advantage. Interestingly, DeepSeek‐R1‐Distill‐Qwen‐1.5B, although distilled from the math‐specialized Qwen‐2.5‐1.5B‐Math model, performs better than its Qwen counterparts, suggesting that the distillation process may have introduced generalization capabilities. The TinyLLaMA variants demonstrate competitive performance, although they generally do not achieve the accuracy of LLaMA-3.2 models. The Pythia‐1B model achieves strong results, in some cases exceeding LLaMA-3.2-1B-Instruct, suggesting it could be a suitable candidate with appropriate training.
@@
Although all models are provided with full context and trained until the trajectory embeddings align with the tokens that enable the LLM to predict delays with minimal error, different LLM backbones still exhibit varying levels of performance. This reflects the importance of the models’ linguistic understanding in achieving accurate post-terminal duration prediction. Thus, it is not only the inclusion of the full context that matters, but also the model’s ability to effectively interpret it to achieve the best predictive results. In addition to assessing the comparative effectiveness of each model, we demonstrate that the proposed framework is compatible with arbitrary LLM backbones by aligning trajectory representations specifically to the selected LLM’s embedding space. This highlights the potential of the proposed cross-modality adaptation technique for application to larger and more capable models.
@@
Our framework achieves high $R^{2}_{D}$ values through a unique problem formulation that leverages information already available to ATCs at the destination aerodrome, namely, the scheduled time of arrival and the actual time the aircraft enters the TMA. This enables efficient prediction without the need for historical delay data or airport network information, unlike the sequential and graph-based models compared in Table 1. By narrowing the focus to a single airport, the framework aligns with ATCs’ operational scope of work and achieves significantly more accurate delay prediction without requiring simultaneous monitoring of multiple airports.
Without a prediction model, ATCs would need to assume an average additional delay of more than 3 minutes per flight at busy airports, as reported in European airport statistics [85], which are comparable to those at Incheon Airport. This could introduce unnecessary overestimation for non-delayed aircraft. In contrast, our framework maintains prediction errors within one minute across several backbones, with LLaMA3.2-1B-Instriuct achieving the best average error of 0.92 minutes. This demonstrates both its accuracy and operational value, indicating its deployability since delays are typically recorded at the minute level [75]. While second-level accuracy may be desirable for certain operations, our framework already meets the current minute-level standard, given the available data precision. Future access to higher-resolution data could further improve accuracy.
@@ 4.3 Effectiveness of Context Inclusion
A key contribution of our proposed framework is the integration of a comprehensive set of information, providing the model with the context of the operational environment. These contextual components are of varying importance; therefore, to evaluate their contributions, we adopt an ablation-based approach in which each component is hierarchically removed from the multimodal scenario representation $s_{i,t}$, followed by retraining and re-evaluation under strictly identical hyperparameter settings. We selected Meta’s LLaMA-3.2-1B-Instruct as the LLM backbone due to its comparatively strong performance, as shown in Table 3. This controlled procedure enables us to directly quantify the performance degradation caused by the absence of each context type, thereby providing empirical evidence of its contribution to the overall predictive capability. Moreover, this analysis was conducted to support our design decision to use full context, demonstrating the value of leveraging multimodal data for accurate delay prediction. The performance of the model under different context removal settings is presented in Table 4 and Figure 7. The importances of context elements are discussed as follows:
@@
  • General Flight Information: The exclusion of basic flight information leads to a performance drop, underscoring the importance of general details of the flight. This finding aligns with prior works, which predominantly rely on general flight information as the primary input, as shown in Table 1.
@@
  • NOTAMs: Excluding NOTAMs shortens the context, which can occasionally reduce the burden of long sequences and improve performance. However, it also removes special operational constraints or disruption, such as runway closures or equipment failures, which can introduce delay [3]. Consequently, performance drops sharply in some months, leading to an overall decline compared to the full context.
@@
  • Trajectory Data: Trajectory data is shown to be the most influential factor for post-terminal duration prediction, as its exclusion results in a severe reduction in performance. This is expected, as trajectory data directly reflects airspace conditions that cannot be inferred from flight or weather information alone. From a technical perspective, removing trajectory inputs also eliminates the learnable portion of the context. This suggests that integrating trajectory data contributes to the predictive accuracy.
@@
    • Focusing Trajectories We find that removing the focusing trajectory leads to a large performance degradation, often comparable to the removal of all trajectory inputs and consistently more severe than excluding other trajectory types. This finding is intuitive and consistent with [25], as the focusing trajectory captures the current flight state under monitoring, while its embedding represents the specific maneuvering procedures most directly related to post-terminal duration.
@@
    • Active Trajectories Excluding the active trajectories from the context removes the model’s awareness of the current airspace conditions, particularly the level of congestion and the overall traffic situation at that time. Therefore, the model cannot account for the congestion factor that often extends the post-terminal duration, making accurate prediction more difficult.
@@
    • Prior Trajectories Since prior trajectories provide patterns that may influence the active and focusing trajectories and indirectly reflect ground traffic, excluding them leaves the model unaware of historical flow patterns in both air and ground operations, thereby decreasing performance.
@@
The experimental results indicate that our accurate predictions are enabled not only by offering a different formulation from prior works but also by incorporating more contextual information within a single airspace. While long contextual inputs can sometimes cause the model to struggle, removing certain elements, such as TAFs and NOTAMs, from many datasets may even improve performance. However, including the full context generally ensures consistent and reliable results, as the model can rely on that information when it is actually required. Since LLMs function as selective sequence encoders that can disregard irrelevant tokens, it is preferable to provide full context and let the model ignore unnecessary tokens rather than risk excluding potentially useful information.
@@
The contribution of textual components follows a similar pattern, though the trajectory components play a more critical role. This is partly because trajectory features are trainable and partly because the post-terminal duration is strongly influenced by prevailing traffic conditions. However, the model still requires the textual components, as removing some or all of them leads to a degradation in performance. The framework demonstrates robustness even when certain contextual information is unavailable, underscoring its suitability for operational use.
@@
Table 4: Average Performance of Meta-LLaMA-3.2-1B-Instruct under Different Context Removal Settings
@@
   Without General Flight Information 0.9738 2.2846 1.4955 11.1673 0.99979 7.8250 0.9142
@@
4.4 Real-time Delay Update
@@
Although the model underestimates the delay in this example, we observe that the absolute error generally remains within one minute, with predictions adapting as the flight progresses. Moreover, the variation in error over time suggests that the model refines its predictions as newly available information updates the prediction context. Importantly, maintaining sub-minute error levels is practically useful for ATM, since delays are typically recorded at the minute level [75]; keeping the error below one minute is therefore considered deployable. Thus, the real-time prediction model from our framework can potentially be integrated into post-terminal ATM systems to assist ATCs in estimating aircraft delays within the terminal area.
@@ 5 Conclusion
This paper presents a flight delay prediction framework that indirectly estimates post-terminal delays by predicting post-terminal duration. The framework is designed to support ATCs’ operational needs by providing accurate, continuously updated delay information. It employs a cross-modality adaptation technique that aligns trajectory data with the LLM by mapping trajectory embeddings from the pre-trained ATSCC encoder to the LLM’s embedding space. The training process remains lightweight, leveraging knowledge from both the trajectory encoder and the LLM, while updating only a cross-modality adaptation network and an output regression head. The framework enables accurate post-terminal duration prediction, with performance gains driven by the comprehensive inclusion of aeronautical information, particularly trajectory data, which allows the model, through the LLM’s linguistic capabilities, to interpret key factors contributing to delays. By incorporating diverse contextual information and focusing prediction within the airspace, our framework achieves sub-minute accuracy and overcomes key limitations of prior works, namely their reliance on prior delay data, fixed structured inputs, and coarse update timescales. In contrast, it focuses on real-time airspace conditions, integrates ATC-accessible information such as scheduled arrival times and airspace entry times, and supports second-level updates for post-terminal ATM operations. Looking forward, the framework can be extended to other phases of delay prediction by incorporating contextual information relevant to those phases. Its generality opens up opportunities for broader ATM tasks, such as trajectory prediction and traffic flow management. Future work may also investigate the integration of higher-quality data and additional modalities, such as images or audio, to further develop multimodal models for aviation applications.